# Estimating Spatially Resolved Radiation Fields Using Neural Networks

### v1

@@ 2.5 Validation Metrics
To assess the quality of the developed models, we need to analyze their performance on the test section of our datasets. As the distribution of voxel values inside our dataset is special for our use case and we have special requirements as well, e.g., we do not aim for perceptual accuracy as in most other reconstruction tasks, we first need to assess which metrics and which variants of metrics are applicable for neural network evaluation of this kind. In general, we measure our metrics over the air kerma rate fields $\dot{K}_{\mathrm{air}}$ of ground-truth and predicted fields using the following three metrics classes.
@@
The first and most obvious kind of metric we employed is the SMAPE which is providing the mean relative error across all voxels of the test dataset. As the dataset is strongly biased towards low voxel values we introduce this metric in two variants: $\mathrm{SMAPE}_{acc}^{x\,\%}$ for the metric applied over all voxels whose values exceed $\max(K_{air})*(1-\frac{x}{100})$ and $\mathrm{SMAPE}_{acc}^{scatter}$, which applies the metric on all voxels between $5\%$ and $0.5\%$ of $\max(K_{air})$.
The SMAPE is defined in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels, present in two corresponding $K_{air}$ fields. $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
@@
This metric is a well established approach of measuring the similarity in terms of structure between a target and a prediction image or volume [40]. This metric is commonly used for reconstruction tasks and expresses the structural similarity between target and prediction by an intuitive value between 0 and $1$, where $1$ indicates perfect accordance between both structures.
@@
The GPR within the scope of a gamma evaluation is a common metric in the field of radiotherapy radiation transport simulation validation [41]. It reflects the relative amount of voxels that pass a certain criterion within a given radiation field. The metric variant we implemented follows the idea of the Maximum Allowed Dose Difference (MADD) as described by Jiang et al.[42]. For our validation we used two different criteria. The strict criteria, which measures the amount of voxels with a MADD of $3\,\%$ using a spatial resolution of $4\,\mathrm{cm}$ and the more relaxed criterion, which measures the amount of voxels with a MADD of $10\,\%$ and a spatial resolution of $6\,\mathrm{cm}$. To calculate the GPR, the field is shifted by the allowed $\Delta_{d}$ in each dimension, respecting the extend of a single voxel. For $\Delta_{d}=4\,\mathrm{cm}$ this results in $2$ shifts per dimension for the case of a voxel size of $2\,\mathrm{cm}$. The minimum distance weighted error for each voxel across the applied shifts is used as the error for the scoring of the passing voxels, thus the error $\gamma(\vec{x})$ for a voxel in a voxel grid $\mathbf{X}$ at position $\vec{x}\in\mathbf{X}$ is defined as:
@@
with $\dot{D}_{t}$ being the target doserate in $K_{air}$ at a given point and $\dot{D}_{p}$ being the predicted one.
@@
As the spectra represent a special kind of information with additional usage aside the direct dosimetry aspect, we evaluate them separately. Therefore, we use the Intersection over Union (IoU), where we calculate the ratio between the combined area of ground truth and prediction spectra and the overlapped area between them. This accuracy is defined as in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels in two corresponding spectra radiation fields. $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
@@ 3 Results
In general, the qualities of the developed models by this research work are condensed in Tab. 3. In there, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets. The first dataset DS-01, was used to train our base FCNN architecture called SRBF. We compared the results of that network against the implementation of the U-Net that we introduced previously as Beam-to-Scatter (B2S) U-Net. As this network does not need any architectural changes to be applied on multiple datasets, as its inputs are the raw primary beam layer of a radiation field together with the X-ray tube output spectra, we used this as a reference for each dataset and variant of our FCNN architecture. We report the best metric values our networks achieved after we optimized the models hyperparameters by using Optuna as mentioned in the previous section. The specific found hyperparameters for each network, are listed in Tab. 4.
On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture. Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharp as the FCNN does. This observation can be reproduced for DS-02. Additionally, we intended to highlight the effectiveness of our spectra encoding for the expected air kerma rates $\dot{K}_{\mathrm{air}}$. Thus, we trained the base SRBFNet model, which can not handle different output spectra of the X-ray tube, on DS-02. The effectivity of our spectra encoding can be seen by comparing SRBFNet against SPERFNet, where we get a nominal difference of $12.9\,\%$ for $\mathrm{SMAPE}_{acc}^{scatter}$ just from the missing spectra information. Furthermore, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution got slightly, but not notably, worse.
For DS-03 and PBRFNet, we observed, that using models with a higher neuron count by increasing the width of our model beyond $192$, e.g., $256$ or even $384$, the reconstruction capability is increasing by up to $1.5\,\%$ regarding $\mathrm{SMAPE}_{acc}^{scatter}$. A similar increase of the SSIM on the other hand side, and thus the overall shape of the radiation field, is invariant to a change of the model width beyond $192$ across all our datasets and models. As the benefit of using broader models is not significant enough to sacrifice performance for it, we limited the hyperparameter selection to models with a maximum model width of $192$.
@@ 4 Discussion
The configuration of the best performing models during the hyperparameter tuning are listed in Tab. 4. From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is the use of sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$. Moreover, when additionally reviewing the table of Tab. 3, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case. Consistently over all datasets, the angular resolution of the beam needed to solve this problem was found to be low enough, that a harmonics count of $4$ for the spherical harmonics encoding of the beam direction is sufficient. Further, two plain FiLM layers for feature fusing, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation. Even though, the impact of the FiLM layers on the inference duration is measurable, it is small compared to the concatenation that is commonly used in the field of Novel View Synthesis, as it can be seen in Tab. 5. In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry. That is, because the staff during an IR procedure will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume. From Tab. 5, one can also draw, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM. Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS. As the provided implementation does not use custom, native GPU codes like implementations in CUDA or OpenCL, its inference time does not enable real-time usage in the sense of the required $8\,\mathrm{ms}$ to $11\,\mathrm{ms}$ for VR- or AR-applications. Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications. After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, like it was already demonstrated by InstantNGP.

### v2

@@ 2.5 Validation Metrics
To assess the quality of the developed models, we need to analyze their performance on the test section of our datasets. As the distribution of voxel values inside our dataset are unique to our use case and we have special requirements as well, we did not aim for perceptual accuracy as in most other reconstruction tasks, we first needed to assess which metrics and which variants of metrics are applicable for neural network evaluation of this kind. In general, we measure our metrics over the air kerma rate free-in-air fields $\dot{K}_{\mathrm{air}}$ of ground-truth and predicted fields using the following three metrics classes.
@@
The first and most obvious kind of metric we employed is the SMAPE which provides the mean relative error across all voxels of the test dataset. As the dataset is strongly biased towards low voxel values we introduced this metric in two variants: $\mathrm{SMAPE}_{acc}^{x\,\%}$ for the metric applied over all voxels whose values exceed $\max(K_{\mathrm{air}})*(1-\frac{x}{100})$ and $\mathrm{SMAPE}_{acc}^{scatter}$, which applies the metric on all voxels between $5\,\%$ and $0.5\,\%$ of $\max(K_{\mathrm{air}})$.
The SMAPE is defined in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels, present in two corresponding $K_{\mathrm{air}}$ fields. $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
@@
This metric is a well established approach to measuring the similarity in terms of structure between a target and a prediction image or volume [40]. This metric is commonly used for reconstruction tasks and expresses the structural similarity between target and prediction by an intuitive value between 0 and $1$, where $1$ indicates perfect accordance between both structures.
@@
The GPR within the scope of a gamma evaluation is a common metric in the field of radiotherapy radiation transport simulation validation [41]. It reflects the relative number of voxels that pass a certain criterion within a given radiation field. The metric variant we implemented follows the idea of the Maximum Allowed Dose Difference (MADD) as described by Jiang et al.[42]. For our validation we used two different criteria. The strict criteria, which measures the number of voxels with a MADD of $3\,\%$ using a spatial resolution of $4\,\mathrm{cm}$ and the more relaxed criterion, which measures the number of voxels with a MADD of $10\,\%$ and a spatial resolution of $6\,\mathrm{cm}$. To calculate the GPR, the field is shifted by the allowed $\Delta_{d}$ in each dimension, respecting the extent of a single voxel. For $\Delta_{d}=4\,\mathrm{cm}$ this results in $2$ shifts per dimension for the case of a voxel size of $2\,\mathrm{cm}$. The minimum distance weighted error for each voxel across the applied shifts is used as the error for the scoring of the passing voxels, thus the error $\gamma(\vec{x})$ for a voxel in a voxel grid $\mathbf{X}$ at position $\vec{x}\in\mathbf{X}$ is defined as:
@@
with $\dot{D}_{t}$ being the target dose rate in $K_{\mathrm{air}}$ at a given point and $\dot{D}_{p}$ being the predicted one.
@@
As the spectra represent a special kind of information that can be used for purposes other than direct dosimetry, we evaluate them separately. Therefore, we use the Intersection over Union (IoU), where we calculate the ratio between the combined area of ground truth and prediction spectra and the overlapped area between them. This accuracy is defined as in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels in two corresponding spectra radiation fields. $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
@@ 3 Results
In general, the qualities of the developed models by this research work are condensed in Tab. 3. Therein, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets. The first dataset DS-01, was used to train our base FCNN architecture called SRBFNet. We compared the results of that network against the implementation of the U-Net that we introduced previously as Beam-to-Scatter (B2S) U-Net. As this network does not need any architectural changes to be applied on multiple datasets, as its inputs are the raw primary beam layer of a radiation field together with the X-ray tube output spectra, we used this as a reference for each dataset and variant of our FCNN architecture. We report the best metric values our networks achieved after we optimized the models hyperparameters by using Optuna as mentioned in the previous section. The specific found hyperparameters for each network, are listed in Tab. 4.
On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture. Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharply as the FCNN does. This observation can be reproduced for DS-02. Additionally, we intended to highlight the effectiveness of our spectra encoding for the expected air kerma rates $\dot{K}_{\mathrm{air}}$. Thus, we trained the base SRBFNet model, which cannot handle different output spectra of the X-ray tube, on DS-02. The effectivity of our spectra encoding can be seen by comparing SRBFNet against SPERFNet, where we get a nominal difference of $12.9\,\%$ for $\mathrm{SMAPE}_{acc}^{scatter}$ just from the missing spectra information. Furthermore, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution got slightly, but not notably, worse.
For DS-03 and PBRFNet, we observed, that using models with a higher neuron count by increasing the width of our model per layer beyond $192$ neurons, e.g., $256$ or even $384$, the reconstruction capability increases by up to $1.5\,\%$ regarding $\mathrm{SMAPE}_{acc}^{scatter}$. A similar increase of the SSIM on the other hand side, and thus the overall shape of the radiation field, is invariant to a change of the model width beyond $192$ across all our datasets and models. As the benefit of using broader models is not significant enough to sacrifice performance for it, we limited the hyperparameter selection to models with a maximum model width of $192$.
@@ 4 Discussion
The configuration of the best performing models during the hyperparameter tuning are listed in Tab. 4. From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is using sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$. Moreover, when additionally reviewing the table of Tab. 3, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case. Consistently over all datasets, the angular resolution of the beam needed to solve this problem was found to be low enough, that a harmonics count of $4$ for the spherical harmonics encoding of the beam direction is sufficient. Further, two plain FiLM layers for feature fusing, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation. Even though, the impact of the FiLM layers on the inference duration is measurable, it is small compared to the concatenation that is commonly used in the field of Novel View Synthesis, as it can be seen in Tab. 5. In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry. That is because the staff, during an IR procedure, will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume. From Tab. 5, one can also draw the conclusion, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM. Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS. As the provided implementation does not use custom, native GPU codes like implementations in CUDA or OpenCL, its inference time does not enable real-time usage in the sense of the required $8\,\mathrm{ms}$ to $11\,\mathrm{ms}$ for VR- or AR-applications. Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications. After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, as has already been demonstrated by InstantNGP.